OpenStudy (anonymous):

what is the equation of a circle with the equation x^2+y^2+2x+4y-9=0 to the nearest thousandth?

OpenStudy (anonymous):

(x+1)^2+(y+2)^2 = 14 x^2+2x+1+y^2+4y+4=9+1+4 add 1 to the x terms above and 4 to the y terms and take the perfect squares of x^2+2x+1 and y^2+4y+4 . make sure and add 1 and 4 to the other side of the equation as well. the circle will have a diameter of /sqrt(14) and a center of (-1, -2)

OpenStudy (anonymous):

not sure what you mean by "to the nearest thousandth" nothing to round there really.